Please help with the homework
At what period will 2500 double if interest is calculated half yearly at 11% interest per annum.
Thanks in advance

Found 2 solutions by Alan3354, ikleyn:
Answer by Alan3354(69443)

At what period will 2500 double if interest is calculated half yearly at 11% interest per annum.
---------------------
11%/year = 5.5% per period
===============
5000 = 2500*(1 + 0.055)^t --- t in periods
1.055^t = 2
t*ln(1.055) = ln(2)
t = ln(2)/ln(1.055)
t = ~ 12.946 periods = 6.47 years

Answer by ikleyn(52787)

.
Please help with the homework
At what period will 2500 double if interest is calculated half yearly at 11% interest per annum.
Thanks in advance
~~~~~~~~~~~~~~~~~~~~~~


            It is  INCORRECTLY  to say   "interest is calculated half yearly at  11%  interest per annum.

            Correctly is to say  "Interest is  COMPOUNDED  every six months at  11%  annual interest rate."



You should calculate the time  "t"  as  Alan did it in his post.


But  AFTER  THAT  you  MUST  to round the time to integer number of compoinding periods,
in order for the bank was in position to make last compounding.


So, in this problem the correct answer is  "6.5  years,  that comprise  13  compounding periods."


//////////////////


To see many other similar  (and different)  solved problems,  look into the lesson
    - Problems on discretely compound accounts
in this site, and learn the subject from there.


After reading this lesson, you will tackle such problems on your own without asking for help from outside.
At the same time,  you will improve your terminology,  which is at low level now.


Also,  you have this free of charge online textbook in ALGEBRA-I in this site
    - ALGEBRA-I - YOUR ONLINE TEXTBOOK.

The referred lesson is the part of this online textbook under the topic "Logarithms".


Save the link to this online textbook together with its description

Free of charge online textbook in ALGEBRA-I
https://www.algebra.com/algebra/homework/quadratic/lessons/ALGEBRA-I-YOUR-ONLINE-TEXTBOOK.lesson

to your archive and use it when it is needed.


Happy learning (!)